OpenStudy (daedae17):

Find the slope of a line parallel to 2x-y=16 A.8 B.-1/2 C.1/2 D.2

OpenStudy (jackellyn):

Lines that are parallel have the same slope. To find the slope of this line, you have to put it in slope-intercept form (y=mx+b). m is the slope. 2x - y = 16 -y = -2x + 16 y = 2x - 16 The slope of this line is 2, and the slope of a parallel line is also 2.

OpenStudy (daedae17):

Thanks ^_^